A car of force 250N acts for 20s.what is the impulse?
1
Expert's answer
2021-09-13T11:01:09-0400

By definition, impulse of a constant force is given as follows:


"J = F\Delta t"

where "F = 250N" is the force that acts during the time "\Delta t = 20s". Thus, obtain:


"J = 250N\cdot 20s = 5000N\cdot s = 5000kg\cdot m/s"

Answer. "5000\space kg\cdot m/s".